a rectangle is 5 times as long as it is wide. If the area is 125 square feet, find its perimeter.
Answer by nyc_function(2741)

length = 5x
width = x
area = 125
A = L*W
125 = (5x)(x)
125 = 5x^2
125/5 = x^2
25 = x^2
sqrt{25} = sqrt{x^2}
5 = x
The width is 5 feet.
The length is 5 times 5 feet or 25 feet.
To find the perimeter, we use P = 2L + 2W.
P = 2(25) + 2(5)
P = 50 + 10
P = 60 feet.
Done!